### v1

# Topological Indices With Degree Sequence $\mathscr{D}$ of Tree

Abstract.
In this paper, we refer to a asymptotic degree sequence as $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. The examination of topological indices on trees gives us a general overview through bounds to find the maximum and minimum bounds which reflect the maximum and minimum number of edges incident to every vertex in the graph, Albertson index known as $\sum_{uv\in E(G)}\lvert d_{u}(G)-d_{v}(G)\rvert$, Sigma index $\sigma(G)$ among $\mathscr{D}$ of tree $T$ when $d_{n}\geqslant\dots\geqslant d_{1}$. According to the first zegrb we show for a degree sequence of order $n=4$, $\operatorname{irr}(T)=M_{1}(T)^{2}-2\sqrt{M_{1}(T)}+\sum_{i=1}^{4}\left|x_{i}-x_{i+1}\right|-(b+c)-1$.
Keywords: Degree, Sequence, Zagreb, Index, Tree.
MSC 2020: 05C05, 05C12, 05C20, 05C25, 05C35, 05C76, 68R10.
MSC 2010: 05C05, 05C12, 05C35, 68R10.
1. Introduction
Throughout this paper, let $G(E,V)$ be a simple, connected graph, denote by $uv$ an edge connecting vertices $u$ and $v$. In 1995, Molloy and Reed in [19] define asymptotic degree sequence as $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$. Zhang et al. in [17] define the set of trees $\mathcal{T}$ among degree sequence $\mathscr{D}$ denote that by $\mathcal{T}_{\mathscr{D}}$, then when $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{n}$, in [21] given $\mathcal{T}_{\mathscr{D}}$ by $\sum_{i=1}^{n}d_{i}=n-1$. The maximum degree $\Delta$ of any tree with degree sequence $\mathscr{D}$ given by Broutin et al. in [20] by $\Delta_{\mathscr{D}}=\max\{i,\mathcal{T}_{\mathscr{D}}>0\}$.
Further, let $\deg i=\deg(i)$ the degree of a vertex $i$, the amount of diverge at $u$ equals $\deg u$, which for any tree has a center chord of a maximum of one or two points. In 1971, let be refer to the maximal number of edges by $\mathcal{E}$, Wai-Kai Chen in [11] (see also [10] refer to the graph $G$ has $n$ nodes and $c$ components, then $\mathcal{E}=\frac{1}{2}(n-c)(n-c+1)$, we can make it as a general case if $G_{i},i=1,2,\ldots,c$, is the components of $G$, each having $n_{i}$ nodes, then $\mathcal{E}=\frac{1}{2}n_{i}\left(n_{i}-1\right)$. In 1997, M. O. Albertson in [12] mention to the imbalance of an edge $uv$ by $imb(uv)=\lvert d_{u}-d_{v}\rvert$, Considered graph $G$ regular if all of its vertices have the exact same degree, then irregularity measure define in [12, 16, 18] as:
$\operatorname{irr}(G)=\sum_{uv\in E(G)}\lvert d_{u}(G)-d_{v}(G)\rvert.$
Dorjsembe et al. In [8] given a relationship with irregularity of Albertson index and minimum, maximum degrees $\delta,\Delta$ of graph $G$, where contribute vital roles in determining connection, shading, component incorporation, and realisability. Among this degrees, which reflect the maximum and minimum number of edges incident to every vertex in the graph, are important elements of a graph’s sophistication and performance. It is given by: $\operatorname{irr}(G)>\frac{\delta(\Delta-\delta)^{2}.|V|}{\Delta+1}$, the degree sum formula (or handshaking lemma) is know well as doubles of edges as $2|E|$. In [14, 18] define total irregularity as $\operatorname{irr}_{T}(G)=\sum_{\{u,v\}\subseteq V(G)}\lvert d_{u}(G)-d_{v}(G)\rvert$ and shown upper bounds on the irregularity of graphs. When $G$ is a tree, Ghalavand, A. et al. In [25] proved $\operatorname{irr}_{T}(G)\leq\frac{n^{2}}{4}\operatorname{irr}(G)$ when the bound is sharp for infinitely many graphs. The first and the second Zagreb index, $M_{1}(G)$ and $M_{2}(G)$ are defined in [13, 15, 12] as:
$M_{1}(G)=\sum_{i=1}^{n}d_{i}^{2},\quad\text{and}\quad M_{2}(G)=\sum_{uv\in E(G)}d_{u}(G)d_{v}(G).$
Actually an alternative expressions introduced by M. Matejí et al. in [12, 2] for the first Zagreb index as $M_{1}(G)=\sum_{u\sim v}(d_{u}+d_{v})$. The recently introduced $\sigma(G)$ irregularity index is a simple diversification of the previously established Albertson irregularity index, in [24, 23] defined as:
$\sigma(G)=\sum_{uv\in E(G)}\left(d_{u}(G)-d_{v}(G)\right)^{2}.$
This paper is organized as follows. In Section 1, we observe the important concepts to our work including literature view of most related papers, in Section 2 We have provided a preface through some of the important theories and properties we have utilised in understanding the work, in Section 3 we introduced the main result of our paper and the goal of this paper about Albertson index among a degree sequence $d_{n}\geqslant\dots\geqslant d_{1}$ of tree, we computing the advanced result of this paper of Sigma index among a degree sequence of tree.
2. Preliminaries
In this section, we considered trees in graphs with fixed number of degree sequence $d_{n}\geqslant\dots\geqslant d_{1}$. In Lemma 2.3, it is explained when a graph is connected if it contains maximal Sigma index consider $\Delta\left(G\right)=\max\left\{{{{\deg}_{G}}(v);v\in V\left(G\right)}\right\}$, that is for any vertex in graph $G$ has a maximum degree of $n-1$, as indicated by the notation $\Delta(G)=n-1$.
Proposition 2.1.
[6] Let be $T$ a tree, a sequence of positive integers ${d_{1}},{d_{2}},{d_{3}};\quad{d_{1}}\geq{d_{2}}\geq{d_{3}}$, Albertson index defined as:
$\operatorname{irr}=\left\{\begin{aligned} &\operatorname{irr}_{max}=(d_{1}-1)^{2}+(d_{2}-1)^{2}+(d_{3}-1)(d_{3}-2)(d_{1}-d_{3})(d_{2}-d_{3})\\ &\operatorname{irr}_{min}=(d_{1}-1)^{2}+(d_{3}-1)^{2}+(d_{2}-1)(d_{2}-2)+(d_{1}-d_{3}).\end{aligned}\right.$
Lemma 2.1.
[1, 4] If $p\geq 1$ is an integer number and a degree sequence $d=(d_{1},d_{2},\ldots,d_{n})$ where $10\leq d\leq n-$ 1, then
$\left(\sum_{i=1}^{n}d_{i}^{p}\right)^{\frac{1}{p}}\leq(n-1)^{1-\frac{1}{p}}\sum_{i=1}^{n}d_{i}^{\frac{1}{p}}$
if and only if $\left(\sum_{i=1}^{n}\sqrt[p]{d_{i}}\right)^{p}\geqslant c\sum_{i=1}^{n}d_{i}^{p}$.
Proposition 2.2.
[3, 18]  Let be a degree sequence $d=(d_{1},\dots,d_{n})$ where $d_{1}\geq d_{2}\geq\dots\geq d_{n}$, then we have
$\sum_{i=1}^{n}d_{i}\leq n(n-1)+\sum_{i=n+1}^{k}\min(n,d_{i}).$
In fact, we employ this proposition in Lemma 3.1 where the degree sequence is $d=(d_{1},d_{2},d_{3},d_{4})$, and $\operatorname{irr}_{t}(G)\leq(n-2)\operatorname{irr}(G)$
Theorem 2.2.
Let be a sequences with $n\geq 3$ and let be order as: ${d_{n}}>d_{1}>\dots>d_{2}>d_{n-1}$, the Caterpillar tree with such order has the maximum value of $\operatorname{irr}$ among all Caterpillar trees with such degrees sequence of path vertices.
Proposition 2.3.
Sigma index for caterpillar tree of order $(n,m)$ given by:
$\sigma(C(n,m))=\begin{cases}2m^{3},&\quad\text{if }n=2\\ 2m^{3}+m-2,&\quad\text{if }n\geqslant 2.\\ \end{cases}$
Proposition 2.4.
For Caterpillar tree with path vertices with degrees $d_{1},d_{2},\dots,d_{n}$, we have:
$\operatorname{irr}(G)=\left({{d_{n}}-1}\right)^{2}+\left({d_{1}-1}\right)^{2}+\sum\limits_{i=2}^{n-1}{\left({{d_{i}}-1}\right)\left({{d_{i}}-2}\right)}+\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|.$
We denote by $\delta(G),\Delta(G)$ the minimum and maximum degrees in a graph $G$, denote by $d(G)$ their average degree. Let $\mathcal{S}$ be a class of graphs, then we have $\operatorname{irr}_{\max},\operatorname{irr}_{\min}$ Albertson index of a graph $G$, where:
$\operatorname{irr}_{\max}(\mathcal{S})=\max\{\operatorname{irr}(G)\mid G\in\mathcal{S}\},$
$\operatorname{irr}_{\min}(\mathcal{S})=\min\{\operatorname{irr}(G)\mid G\in\mathcal{S}\}.$
Lemma 2.3 (Yang J, Deng H., M., [7]).
Let $G$ has the maximal Sigma index among all connected graphs with $n$ vertices and $p$ pendant vertices, where $n,p$ are positive integers such that $31\leq p\leq n-$ 3. Then: $\Delta(G)=n-1$.
Oboudi, M., R. in [9] shows if $G$ is a connected graph then we have a sequence $d_{1}>d_{2}>\dots>d_{n}$ that has $n$ degree graph and $n\geq 3$. The first and second Zagreb indices in [24] as: Two additional Zagreb-type indices are the forgotten index in [24] as:
$F(G)=\sum_{w\in E(G)}\left(d_{w}\right)^{3}=\sum_{uv\in E(G)}\left[\left(d_{u}\right)^{2}+\left(d_{v}\right)^{2}\right].$
Theorem 2.4.
[26] Let be a complete bipartite graph is $K_{m,n}$, then sigma index of graph is $\sigma(K_{m,n})=(mn)(n-m)^{2}$.
Theorem 2.5 (Sigma index  ).
[24] For every simple graph $G,\sigma(G)$ is an even integer, so that we have:
$\sigma(G)=F(G)-2M_{2}(G)=\sum_{uv\in E(G)}\left[\left(d_{u}\right)^{2}+\left(d_{v}\right)^{2}\right]-2\sum_{uv\in\in(G)}d_{u}d_{v}.$
where $M_{2}(G)=\sum_{uv\in E(G)}d_{u}d_{v}$.
Theorem 2.6.
[24] For any double star graph $S_{r,k}$, the degrees of two adjacent central vertices $u,v$ defined as $d_{u}=k\geq 3,d_{v}=r\geq 1$. Then the sigma index of $S_{r,k}$ is given by
$\sigma\left(S_{r,k}\right)=(k-1)^{3}+k^{2}+(r-1)^{3}+r^{2}-2kr.$
Proposition 2.5.
Let be $\mathscr{G}(n)$ is the set of all $n$ vertex on graph $G$, $\lambda(G)$ the index of the largest eigenvalue in graph, and let be $\bar{d}(G)$ the mean of the vertex degrees, then we have:
$\max\{\operatorname{irr}(G):G\in\mathscr{G}(n)\}=\left\{\begin{array}[]{ll}\frac{1}{4}n-\frac{1}{2}&(n\text{ even })\\ \frac{1}{4}n-\frac{1}{2}+\frac{1}{4n}&(n\text{ odd }).\end{array}\right.$
This result implies in [29, 28, 12, 16]. This relationship can be demonstrated for the indices in both Lemma 2.7, 2.8, which depict the star graph.
Lemma 2.7.
[27]  The star $S_{n}$ is the only tree of order n that has a great deal of irregularity, satisfying:
$\operatorname{irr}\left({{S_{n}}}\right)=\left({n-2}\right)\left({n-1}\right)$
Lemma 2.8.
Let be $\mathcal{T}$ a classes of trees, then $\sigma_{max},\sigma_{min}$ in trees with $n$ vertices by:
$\sigma(\mathcal{T})=\left\{\begin{array}[]{l}{\sigma_{\max}(\mathcal{T})}=\left({n-1}\right)\left({n-2}\right){\rm{}};n\geq 3\\ {\sigma_{\min}(\mathcal{T})}=0{\rm{;n=2}}\end{array}\right.$
According to a tree with $n$ vertices has $n-1$ edges, then we have $\sigma_{\max}(T)=n-2$. Thus, the modified total Sigma as $\sigma_{t}(T)=\frac{1}{2}\sum_{(u,v)\subseteq V(T)}\left(d_{u}-d_{v}\right)^{2}$
3. Main Result
In this section, we provide the main result in this paper, we show that in Theorem 3.6 by study properties in Lemma 3.2,3.5. We mention by $C(n,m)$ caterpillar tree with $(n,m)$ vertices, and refer to $n$ main vertices, we have $X=(x_{1},x_{2},\dots,x_{n})$ where $n=|X|$, and refer to $m$ a set of leaves adjacent to main vertices. Zhang, P., Wang, X in[5] show “Caterpillar trees” are employed in chemical graph theory for describing molecular structures, and several topological indices, which include the Albertson index, have been used to study the characteristics of molecules.
3.1. Albertson Index With Degree Sequence $d_{n}\geqslant\dots\geqslant d_{1}$
Hypothesize 1.
Let be a sequences with $n\geq 3$ and let be order as: ${d_{n}}>d_{n-1}>\dots>d_{2}>d_{1}$, the Caterpillar tree with such order has the minimum value of $\operatorname{irr}$ among all Caterpillar trees with such degrees sequence of path vertices.
Proof.
Let be assume the sequence is $d=(d_{1},d_{2},\dots,d_{n})$ where ${d_{n}}>d_{n-1}>\dots>d_{2}>d_{1}$, then we have:
$\operatorname{irr}(G)$ $=\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}(d_{i}-1)(d_{i}-2)+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}$
$=\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+2(n-3)+(d_{n}^{2}-2d_{n}+1)+(d_{n-1}^{2}-2d_{n-1}+1)$
$=\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}d_{i}^{2}+d_{n}^{2}+d_{n-1}^{2}-3\sum_{i=2}^{n-2}d_{i}-2d_{n}-2d_{n-1}+2(n-3)+2$
$=\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}d_{i}^{2}+d_{n}^{2}+d_{n-1}^{2}-3\sum_{i=2}^{n-2}d_{i}-2d_{n}-2d_{n-1}+2n-4$
$=\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}+d_{n-1}^{2}-2(d_{n}+d_{n-1})+2n-4.$
where we have $d_{n}+d_{n-1}$ is the minimum when ${d_{n}}>d_{n-1}>\dots>d_{2}>d_{1}$, then we have the cases are:
Case 1:
we know for $\sum_{i=2}^{n-2}d_{i}^{2},3\sum_{i=2}^{n-2}d_{i}$, if $d_{i}$ are all 1, then $\sum d_{i}^{2}=n-3$ and $3\sum d_{i}=3(n-3)$, so in this case, $(d_{i})^{2}$ is the minimum. Also, we have from this inequity $d_{n}^{2}+d_{n-1}^{2}$ is still minimum when ${d_{n}}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$.
Case 2:
For $(d_{i})^{2},|d_{i}-d_{i+1}|$, we know if $d_{i}>d_{i+1}$, then $d_{i}^{2}\geq(d_{i}-d_{i+1})$, then we have:
$\operatorname{irr}(G)$ $=(d_{1}-d_{n})+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}+2n-6$
$=d_{1}+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n-1}^{2}-2d_{n-1}+1+d_{n}^{2}-2d_{n}+1+2n-6$
$=d_{1}+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}+d_{n-1}^{2}-2(d_{n}+d_{n-1})+2n-4.$
Let be the sequence $d=(d_{1},\dots,d_{n})$ where $d_{n}>d_{1}>d_{n-2}>d_{3}\dots>d_{2}>d_{n-1}$ or If $d_{i}\leq d_{i+1}$, then we have $d_{i+1}$ as $d_{i}+k$ where $k>0$. where $\min\{d_{n},d_{1}\}=d_{n},(d_{n}-1)^{2}=d_{n}^{2}-2d_{n}+1$, thus $(d_{i})^{2}>|d_{i}-d_{i+1}|$. Therefore we have:
$\operatorname{irr}(G)$ $=(d_{n}-d_{1})+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}+2n-6$
$=-d_{1}+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}+d_{n-1}^{2}-2(d_{n}+d_{n-1})+2n-4.$
Thus, from both the previously discussed cases 1 and 2, we observed that the relation is true for $n$ and therefore the relation has the minimum value of the Albertson index when ${d_{n}}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$ and now we need to prove it for $n+1$ where we consider both $\sum_{i=2}^{n-2}(d_{i}-1)(d_{i}-2)$ as constants and $d_{n}+d_{n-1}$ as the minimum value and to prove that the relation $\sum_{i=2}^{n}|d_{i}-d_{i+1}|$ is as minimum as possible for the Albertson index according to the following:
$\operatorname{irr}(G)$ $=\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}(d_{i}-1)(d_{i}-2)+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}$
$=\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}+d_{n-1}^{2}-2(d_{n}+d_{n-1})+2n-4.$
Clarity, we notice $\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}+d_{n-1}^{2}-2(d_{n}+d_{n-1})+2n-4$ is the minimum when ${d_{n}}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$, then we have:
  • If $d_{i}\geq d_{i+1}$, then we have $\operatorname{irr}(G)=d_{1}+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}-3d_{n}+d_{n-1}^{2}-3d_{n-1}+2n-5$.
  • If $d_{i}\leq d_{i+1}$, then we have: $\operatorname{irr}(G)=-d_{1}+\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n}^{2}-d_{n}+d_{n-1}^{2}-d_{n-1}+2n-5$.
From this cases notice $\sum_{i=2}^{n-2}d_{i}^{2}-3\sum_{i=2}^{n-2}d_{i}+d_{n-1}^{2}+d_{n}^{2}+2n-5$ is minimum according to the constant term, then $\sum_{i=1}^{n-1}|d_{i}-d_{i+1}|+\sum_{i=2}^{n-2}(d_{i}-1)(d_{i}-2)+(d_{n}-1)^{2}+(d_{n-1}-1)^{2}$ is minimum when ${d_{n}}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$. As desire. ∎
Lemma 3.1.
Let be $T$ a tree with a degree sequence of positive integers vertices $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{1}\geq d_{2}\geq d_{3}\geq d_{4}$ we define for numbers $a,b,c,d$ that:
$\operatorname{irr}(T)=M_{1}(T)^{2}-2\sqrt{M_{1}(T)}+\sum_{i=1}^{4}\left|x_{i}-x_{i+1}\right|-(b+c)-1.$
Proof.
By using Albertson index we define the sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{1}\geq d_{2}\geq d_{3}\geq d_{4}$, then by considering $M_{1}(G)=\sum_{i=1}^{n}d_{i}^{2}$ we discus according to Proposition 2.2 the following cases.
Case 1:
In this case, we have the largest value $a,d$, the smallest value $b,c$, then we have:
$\operatorname{irr}(T_{1})$ $=(a-b)+|b-c|+(d-c)-(b+c)$
$=(a+d)-(b+c)+|b-c|-(b+c)$
$=(a+d)-2(b+c)+|b-c|$
$=|a-b|+|b-c|+|c-d|-(b+c)\quad\text{where }|b-c|=-2d_{3}-2d_{4}$
$=\left(d_{1}+d_{2}\right)-2\left(d_{3}+d_{4}\right)$
$=d_{1}+d_{2}-2d_{3}-2d_{4}.$
Case 2:
In this case, we have the largest value $b,c$, the smallest value $a,d$, then we have:
$\operatorname{irr}(T_{2})$ $=b-a+|b-c|+c-d-(b+c)$
$=B-L-B+|b-c|=|b-c|-L\quad\text{where }B=\sum(a,c),L=\sum(a,d),|b-c|=d_{1}-d_{2}$
$=d_{1}-d_{2}-d_{3}-d_{4}.$
We noticed that from both cases $\operatorname{irr}(T_{1})+\operatorname{irr}(T_{2})<\operatorname{irr}(T_{1})-\operatorname{irr}(T_{2})$, then we have $\max(\operatorname{irr}(T_{1}),\operatorname{irr}(T_{2}))=2d_{2}-d_{3}-d_{4}$ and $\min(\operatorname{irr}(T_{1}),\operatorname{irr}(T_{2}))=2d_{1}-3d_{3}-3d_{4}$.
Case 3:
In this case, the largest value is $a,b$, the smallest value is $d,c$, then we have:
$\operatorname{irr}(T_{3})$ $=|a-b|+b-c+|c-d|-b-c$
$=d_{1}-d_{2}+d_{3}-d_{4}-2C\quad\text{where }C\in\left\{d_{3},d_{4}\right\}$
$=d_{1}-d_{2}-d_{3}-d_{4}\quad\text{ if }c=d_{3}$
$=d_{1}-d_{2}+d_{3}-3d_{4}\quad\text{ if }c=d_{4}.$
Case 4:
In this case, the largest value is $a,c$, the smallest value is $d,b$, then we have:
$\operatorname{irr}(T_{4})$ $=a-b+c-b+c-d-b-c$
$=a-3b+c-d$
$=d_{1}-d_{2}-3d_{3}-d_{4}\quad\text{ if }b=d_{3}$
$=d_{1}+d_{2}-d_{3}-3d_{4}\quad\text{ if }b=d_{4}.$
Therefore, from cases $\operatorname{irr}(T_{1}),\operatorname{irr}(T_{2}),\operatorname{irr}(T_{3}),\operatorname{irr}(T_{4})$ we have after comparing
$\operatorname{irr}(T)=\begin{cases}d_{1}+d_{2}-d_{3}-3d_{4}:\quad\text{ the biggest }\\ d_{1}-d_{2}-d_{3}-d_{4}:\quad\text{ the smallest }.\end{cases}$
We should be know: $a\in\left\{d_{1},d_{2}\right\},b\in\left\{d_{3},d_{4}\right\},c\in\left\{d_{3},d_{4}\right\},d\in\left\{d_{1},d_{2}\right\}$
So that we have: $2d_{2}-d_{3}\geq 0$; so, we obtain on: If $a,d\in\left\{d_{1},d_{2}\right\}$ and $b,c\in\left\{d_{3},d_{4}\right\}$ or $a,c\in\left\{d_{1},d_{2}\right\}$ and $b=d_{4},d=d_{3}$, we have:
$\operatorname{irr}_{\max}(T)=\sum_{i=1}^{4}\left(d_{i}-1\right)^{2}+d_{1}+d_{2}-d_{3}-3d_{4}+2.$
And we have also: If $a,d\in\left\{d_{3},d_{4}\right\}$ and $b,c\in\left\{d_{1},d_{2}\right\}$ or $a,\mathrm{~{}b}\in\left\{d_{1},d_{2}\right\}$ and $\mathrm{d}=d_{4},\mathrm{c}=d_{3}$, we have
$\operatorname{irr}_{\min}(T)=\sum_{i=1}^{4}\left(d_{i}-1\right)^{2}+d_{1}-d_{2}-d_{3}-d_{4}+2.$
Therefore, we have
$\operatorname{irr}(G)$ $=(a-1)^{2}+(b-2)(b-1)+(c-2)(c-1)+(d-1)^{2}+|a-b|+|b-c|+|c-d|$
$=(a-1)^{2}+(b-1)^{2}-(b-1)+(c-1)^{2}-(c-1)+(d-1)^{2}+(a-b)+(b-c)+|c-d|$
$=\sum_{x\in\{a,b,d\}}(x-1)^{2}+|a-b|+|b-c|+|c-d|-(b+c)+2$
Finally, according to the largest value and smallest value, we organized that as the maximum value and minimum value, thus we have:
$\operatorname{irr}(T)=\begin{cases}&\operatorname{irr}_{\max}(T)=\sum_{i=1}^{4}\left(d_{i}-1\right)^{2}+d_{1}+d_{2}-d_{3}-3d_{4}+2\\ &\operatorname{irr}_{\min}(T)=\sum_{i=1}^{4}\left(d_{i}-1\right)^{2}+d_{1}-d_{2}-d_{3}-d_{4}+2.\end{cases}$
As desired. ∎
Lemma 3.2.
Let be $T$ a tree of order $n\geqslant 4$, define the sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Albertson index among tree $T$ is:
$\operatorname{irr}(T)=\begin{cases}\operatorname{irr}_{\max}(T)=d_{1}^{2}+d_{2}^{2}+\sum_{i=1}^{3}\lvert d_{i}-d_{i+2}\rvert+d_{3}^{2}+d_{4}^{2}+d_{3}+d_{4}-6\\ \operatorname{irr}_{\min}(T)=d_{3}^{2}+d_{4}^{2}+\sum_{i=1}^{2}\lvert d_{i}-d_{i+2}\rvert+\lvert d_{1}-d_{2}\rvert+d_{1}^{2}+d_{2}^{2}+d_{1}+d_{2}-6.\end{cases}$
Lemma 3.3.
Let be $T$ a tree with a degree sequence of positive integers $d_{i}=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{5}$ and $d_{j}=(d_{1},\dots,d_{6})$ where $d_{1}\geqslant\dots\geqslant d_{n}$ where $i,j>0$ and let be define positive numbers $a,b,c,d,e$, Albertson index among tree $T$ is:
$\operatorname{irr}(T)=\sum_{i=1}^{5}(d_{i}-1)(d_{i}-2)+\sum_{i=1}^{5}|d_{i}-d_{i+1}|+(d_{1}-1)^{2}+(d_{5}-1)^{2}.$
and
$\operatorname{irr}(T)=\sum_{i=1}^{6}(d_{i}-1)(d_{i}-2)+\sum_{i=1}^{6}|d_{i}-d_{i+1}|+(d_{1}-1)^{2}+(d_{6}-1)^{2}.$
Lemma 3.4.
Let be $T$ tree of order $n\geqslant 5$, a degree sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Albertson index among tree $T$ is:
$\operatorname{irr}(T)=d_{1}^{2}+d_{n}^{2}+\sum_{i=2}^{n-1}\lvert d_{i}-d_{i+1}\rvert+\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)-2.$
Proof.
According to Lemma 3.2, we will employed that for a degree sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, therefore we will discus following cases by considering the constant term is $\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)$.
Case 1:
In this case, the largest value is $d_{1},d_{5}$, the smallest value is $d_{2},d_{3},d_{4}$, then we have following cases.
  • In this case, the sequence is $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$, then we have
    $\operatorname{irr}(T_{1})$ $=(d_{1}+1)(d_{1}-1)+(d_{5}+1)(d_{5}-1)+\lvert d_{1}-d_{2}\rvert+\lvert d_{2}-d_{3}\rvert+\lvert d_{3}-d_{4}\rvert+\lvert d_{4}-d_{5}\rvert+$
    $+(d_{2}+2)(d_{2}-1)+(d_{3}+2)(d_{3}-1)+(d_{4}+2)(d_{4}-1)$
    $=d_{1}^{2}+d_{5}^{2}+\sum_{i=1}^{4}\lvert d_{i}-d_{i+1}\rvert+\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)-2.$
  • In this case, the sequence is $d=(d_{1},d_{3},d_{2},d_{4},d_{5})$, the constant term is $\lvert d_{3}-d_{2}\rvert+\lvert d_{4}-d_{5}\rvert-2$, then we have,
    $\operatorname{irr}(T_{2})$ $=(d_{1}+1)(d_{1}-1)+(d_{5}+1)(d_{5}-1)+\lvert d_{1}-d_{3}\rvert+\lvert d_{3}-d_{2}\rvert+\lvert d_{2}-d_{4}\rvert+\lvert d_{4}-d_{5}\rvert+$
    $+\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)$
    $=d_{1}^{2}+d_{5}^{2}+\sum_{i\in\{1,2\}}\lvert d_{i}-d_{i+2}\rvert+\lvert d_{3}-d_{2}\rvert+\lvert d_{4}-d_{5}\rvert+\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)-2.$
  • In this case, the sequence is $d=(d_{1},d_{3},d_{4},d_{2},d_{5})$, the constant term is $\lvert d_{3}-d_{4}\rvert+\lvert d_{2}-d_{5}\rvert-2$, then we have,
    $\operatorname{irr}(T_{3})$ $=(d_{1}+1)(d_{1}-1)+(d_{5}+1)(d_{5}-1)+\lvert d_{1}-d_{3}\rvert+\lvert d_{3}-d_{4}\rvert+\lvert d_{4}-d_{2}\rvert+\lvert d_{2}-d_{5}\rvert+$
    $+\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)$
    $=d_{1}^{2}+d_{5}^{2}+\sum_{i\in\{1,2\}}\lvert d_{i}-d_{i+2}\rvert+\lvert d_{3}-d_{4}\rvert+\lvert d_{2}-d_{5}\rvert+\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)-2$
    $=d_{1}^{2}+d_{5}^{2}+\sum_{i\in\{1,2\}}\lvert d_{i}-d_{i+2}\rvert-d_{1}+2d_{4}-2d_{2}+d_{5}+\sum_{i=2}^{4}(d_{i}+2)(d_{i}-1)-2.$
Case 2:
In this case, the largest value is $d_{2}$, the smallest value is $d_{1},d_{3},d_{4},d_{5}$, then we have following cases.
  • In this case, the sequence is $d=(d_{2},d_{3},d_{1},d_{5},d_{4})$, then we have
    $\operatorname{irr}(T_{4})$ $=d_{2}^{2}+d_{4}^{2}+|d_{2}-d_{3}|+|d_{3}-d_{1}|+|d_{1}-d_{5}|+|d_{5}-d_{4}|+\sum_{i\in\{1,3,5\}}(d_{i}+2)(d_{i}-1)-2$
    $=d_{2}^{2}+d_{4}^{2}+2d_{5}+2d_{3}-2d_{1}-d_{2}-d_{4}+\sum_{i\in\{1,3,5\}}(d_{i}+2)(d_{i}-1)-2.$
  • In this case, the sequence is $d=(d_{2},d_{5},d_{3},d_{1},d_{4})$, then we have
    $\operatorname{irr}(T_{5})$ $=d_{2}^{2}+d_{4}^{2}+|d_{2}-d_{5}|+|d_{5}-d_{3}|+|d_{3}-d_{1}|+|d_{1}-d_{4}|+\sum_{i\in\{1,3,5\}}(d_{i}+2)(d_{i}-1)-2$
    $=d_{2}^{2}+d_{4}^{2}+2d_{5}-2d_{1}-d_{2}+d_{4}+\sum_{i\in\{1,3,5\}}(d_{i}+2)(d_{i}-1)-2.$
  • In this case, the sequence is $d=(d_{2},d_{1},d_{5},d_{3},d_{4})$, then we have
    $\operatorname{irr}(T_{6})$ $=d_{2}^{2}+d_{4}^{2}+|d_{2}-d_{1}|+|d_{1}-d_{5}|+|d_{5}-d_{3}|+|d_{3}-d_{4}|+\sum_{i\in\{1,3,5\}}(d_{i}+2)(d_{i}-1)-2$
    $=d_{2}^{2}+d_{4}^{2}+2d_{5}-2d_{1}+d_{2}-2d_{3}+d_{4}+\sum_{i\in\{1,3,5\}}(d_{i}+2)(d_{i}-1)-2.$
Case 3:
In this case, the largest value is $d_{3},d_{5}$, the smallest value is $d_{1},d_{2},d_{4}$, then we have following cases.
  • In this case, given the sequence as $d=(d_{3},d_{1},d_{2},d_{4},d_{5})$, then we have
    $\operatorname{irr}(T_{7})$ $=d_{3}^{2}+d_{5}^{2}-2d_{1}+d_{3}+d_{5}+\sum_{i\in\{1,2,4\}}(d_{i}+2)(d_{i}-1)-2$
    $=d_{3}^{2}+d_{5}^{2}+\sum_{i\in\{1,2\}}\lvert d_{i}-d_{i+2}\rvert+\sum_{i\in\{1,4\}}\lvert d_{i}-d_{i+1}\rvert+\sum_{i\in\{1,2,4\}}(d_{i}+2)(d_{i}-1)-2.$
  • In this case, given the sequence as $d=(d_{3},d_{2},d_{4},d_{1},d_{5})$, then we have
    $\operatorname{irr}(T_{8})$ $=d_{3}^{2}+d_{5}^{2}-2d_{1}-2d_{2}+d_{3}+2d_{4}+d_{5}+\sum_{i\in\{1,2,4\}}(d_{i}+2)(d_{i}-1)-2$
    $=d_{3}^{2}+d_{5}^{2}+|d_{3}-d_{2}|+|d_{2}-d_{4}|+|d_{4}-d_{1}|+|d_{1}-d_{5}|+\sum_{i\in\{1,2,4\}}(d_{i}+2)(d_{i}-1)-2.$
  • In this case, the sequence is $d=(d_{3},d_{4},d_{1},d_{2},d_{5})$, we notice that in this case the constant term is $|d_{4}-d_{1}|+|d_{2}-d_{5}|-2$, then we have
    $\operatorname{irr}(T_{9})$ $=d_{3}^{2}+d_{5}^{2}+|d_{3}-d_{4}|+|d_{4}-d_{1}|+|d_{1}-d_{2}|+|d_{2}-d_{5}|+\sum_{i\in\{1,2,4\}}(d_{i}+2)(d_{i}-1)-2$
    $=d_{3}^{2}+d_{5}^{2}+\sum_{i\in\{1,3\}}\lvert d_{i}-d_{i+1}\rvert+|d_{4}-d_{1}|+|d_{2}-d_{5}|+\sum_{i\in\{1,2,4\}}(d_{i}+2)(d_{i}-1)-2.$
Actually, respecting the constant terms in each case that shift with the shift of vertices, we notice in Case 1 for $\operatorname{irr}(T_{2}),\operatorname{irr}(T_{3})$, the maximum value is $\operatorname{irr}(T_{3})$ and the minimum value is $\operatorname{irr}(T_{2})$, is available for each case, another aspect to note is that there are repeated results by comparing the vertices with each other, so we restrict ourselves to these comparisons, which they have distributed through three main cases that give us the maximum and minimum values in this sequence, considering the constant terms in each of them.
Thus, we have:
$\operatorname{irr}(T)=\begin{cases}\operatorname{irr}_{\max}(T)=d_{3}^{2}+d_{5}^{2}-2d_{1}-2d_{2}+d_{3}+2d_{4}+d_{5}+\sum_{i\in\{1,2,4\}}(d_{i}+2)(d_{i}-1)-2\\ \operatorname{irr}_{\min}(T)=d_{2}^{2}+d_{4}^{2}+2d_{5}+2d_{3}-2d_{1}-d_{2}-d_{4}+\sum_{i\in\{1,3,5\}}(d_{i}+2)(d_{i}-1)-2.\end{cases}$
As desire. ∎
Lemma 3.5.
Let be $T$ is a tree of order $n$, degree sequence $d=(d_{1},d_{2},\dots,d_{6})$ where $d_{6}\geqslant\dots\geqslant d_{1}$, then Albertson index over tree $T$ is given by:
$\operatorname{irr}(T)=\begin{cases}\operatorname{irr}_{\max}(T)=d_{1}^{2}+d_{6}^{2}+\sum_{i=1}^{6}\lvert d_{i}-d_{i+1}\rvert+\sum_{i=2}^{5}(d_{i}+2)(d_{i}-1)-2\\ \operatorname{irr}_{\min}(T)=d_{1}^{2}+d_{2}^{2}+\sum_{i=1}^{6}\lvert d_{i}-d_{i+1}\rvert+\sum_{i\in\{3,4,5,6\}}(d_{i}+2)(d_{i}-1)-2.\end{cases}$
According to the results we gave in both Lemmas 3.5, 3.4 and 3.2, we employ them in Theorem 3.6 for degree sequence is $d=(d_{1},\dots,d_{n})$ where $d_{n}\geqslant\dots\geqslant d_{1}$ as we show that.
Theorem 3.6.
Let be $T$ tree of order $n$, a degree sequence is $d=(d_{1},\dots,d_{n})$ where $d_{n}\geqslant\dots\geqslant d_{1}$, then Albertson index of tree $T$ is:
$\operatorname{irr}(T)=d_{1}^{2}+d_{n}^{2}+\sum_{i=2}^{n-1}d_{i}^{2}+\sum_{i=2}^{n-1}d_{i}+d_{n}-d_{1}-2n+2.$
Proof.
Let be assume the sequence is $d=(d_{1},\dots,d_{n})$ where $d_{n}\geqslant\dots\geqslant d_{1}$, then by considering measures the total absolute difference between consecutive terms are $\sum_{i=1}^{n-1}(d_{i+1}-d_{i})=(d_{2}-d_{1})+(d_{3}-d_{2})+\dots+(d_{n}-d_{n-1})=d_{n}-d_{1}$. In performing the proof of the relationship, we use mathematical induction, where we find that the relationship is clearly true for certain values, so we will prove the validity of the relationship for $n$ and then for $n+1$.
Case 1:
Assume the relationship is correct for some value and we need to prove that for $n$ where $n\geq 3$, then we have
$\operatorname{irr}(T)$ $=\sum_{i=1}^{n}\lvert d_{i}-d_{i+1}\rvert+(d_{1}^{2}-1)+(d_{n}^{2}-1)+\sum_{i=2}^{n-1}\lvert d_{i}+2\rvert\lvert d_{i}-1\rvert$
$=(d_{n}-d_{1})+(d_{1}^{2}-1)+(d_{n}^{2}-1)+\sum_{i=2}^{n-1}(d_{i}+2)(d_{i}-1)$
$=d_{1}^{2}+d_{n}^{2}+\sum_{i=2}^{n-1}d_{i}^{2}+\sum_{i=2}^{n-1}d_{i}+d_{n}-d_{1}-2(n-2)-2.$
It is clear to show $\sum_{i=2}^{n-1}d_{i}^{2}+\sum_{i=2}^{n-1}d_{i}>d_{1}^{2}+d_{n}^{2}$ by considering the constant term is $d_{n}-d_{1}-2n+2$, in this case, there are multiple approaches that tackle irregularity have certain of those measurements are in concepts, then by considered the quantity $\sum_{i=2}^{n-1}d_{i}^{2}+\sum_{i=2}^{n-1}d_{i}\leqslant(n-1)(n-1)^{2}+2$, So that the relationship is correct for $n$ when $d_{n}\geqslant\dots\geqslant d_{1}$.
Case 2:
Now we need to prove that for $n+1$, by considering $\sum_{i=2}^{n}(d_{i+1}-d_{i})=(d_{3}-d_{2})+(d_{4}-d_{3})+\dots+(d_{n+1}-d_{n})=d_{n+1}-d_{2}$, The expression holds for $n\geq 3$, then we have
$\operatorname{irr}(T)$ $=\sum_{i=2}^{n+1}\lvert d_{i}-d_{i+1}\rvert+(d_{2}^{2}-1)+(d_{n+1}^{2}-1)+\sum_{i=3}^{n}\lvert d_{i}+2\rvert\lvert d_{i}-1\rvert$
$=(d_{n+1}-d_{2})+(d_{2}^{2}-1)+(d_{n+1}^{2}-1)+\sum_{i=3}^{n}|d_{i}+2||d_{i}-1|$
$=d_{n+1}^{2}+d_{2}^{2}+\sum_{i=3}^{n}d_{i}^{2}+\sum_{i=3}^{n}d_{i}+d_{n+1}-d_{2}-2(n-2).$
Therefore the constant term is $(d_{n+1}-d_{2})+(d_{2}^{2}-1)+(d_{n+1}^{2}-1)$ and $(d_{n}-d_{1})+(d_{1}^{2}-1)+(d_{n}^{2}-1)$ it is hold to same result in Case 1, thus the relation is correct for $n+1$. So from both Cases 1 and 2, we find that the relationship is realised.
As desire. ∎
3.2. Sigma Index With Degree Sequence $d_{n}\geqslant\dots\geqslant d_{1}$
Lemma 3.7.
Let be $T$ tree of order $n\leqslant 3$, a sequence $d=(d_{1},d_{2},d_{3})$ where $d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index is:
$\sigma(T)=\begin{cases}\sigma_{\max}(T)=\sum_{i\in\{2,3\}}(d_{i}+1)(d_{i}-1)^{2}+(d_{3}-d_{1})^{2}+(d_{1}-d_{2})^{2}\\ \sigma_{\min}(T)=\sum_{i\in\{1,3\}}(d_{i}+1)(d_{i}-1)^{2}+(d_{1}-d_{3})^{2}+(d_{3}-d_{2})^{2}.\end{cases}$
Furthermore, in Lemma 3.8 we provide a sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{4}\geqslant\dots\geqslant d_{1}$ to define Sigma index among tree with determine sequence as we show that in Lemma 3.7 where the vertices is located as random and by respectively the largest and smallest vertex.
Lemma 3.8.
Let $T$ be a tree of order $n>0$ and a sequence $d=(d_{1},d_{2},d_{3},d_{4})$ where $d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index is:
$\sigma(T)=\begin{cases}\sigma_{\max}(T)=\sum_{i\in\{2,3\}}(d_{i}+1)(d_{i}-1)^{2}+\sum_{i\in\{1,4\}}(d_{i}+2)(d_{i}-1)^{2}+\sum_{i=1}^{4}(d_{i}-d_{i+2})^{2}+(d_{4}-d_{1})^{2}\\ \sigma_{\min}(T)=\sum_{i\in\{1,4\}}(d_{i}+1)(d_{i}-1)^{2}+\sum_{i=1}^{4}(d_{i}-d_{i+2})^{2}+\sum_{i\in\{2,3\}}(d_{i}+2)(d_{i}-1)^{2}.\end{cases}$
According to Lemma 3.8,3.7, in Hypothesize 2, we present a preliminaries for the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$ in Lemma LABEL:le.sigma5 Sigma index for determine sequence.
Hypothesize 2.
Let $T$ tree of order $n>0$, and let be the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$, then Sigma index of tree is:
$\sigma(T)=\sum_{i=1}^{4}d_{i}(d_{i+1})^{2}+(d_{1}-1)^{3}+(d_{4})^{3}+\sum_{i=1}^{5}(d_{i}-d_{i+1})^{2}.$
where inequality holds if and only if $d_{i}=d_{i-1}+1$.
Proof.
According to the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5})$ where $d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geq d_{1}$, we have many cases of position the vertices in tree $T$. In fact, the order in which the elements of this sequence can be realised is 120, and in order to choose the maximum and minimum cases, we get 10 cases, which we can demonstrate as follows:
Case 1:
In this case, we will consider the largest is $d_{1},d_{5}$ by employee $d_{1}$ as constant term, then we have:
$\sigma(T_{1})=\begin{cases}3d_{1}^{3}-2d_{1}^{2}+4d_{1}+d_{2}d_{3}^{2}+d_{4}^{3}+3&d=(d_{1},d_{2},d_{3},d_{4},d_{5}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{2}^{3}+3&d=(d_{1},d_{2},d_{3},d_{5},d_{4}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{3}^{3}+14&d=(d_{1},d_{2},d_{5},d_{3},d_{4}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{3}^{3}+19&d=(d_{1},d_{5},d_{2},d_{3},d_{4}),\\ 3d_{1}^{3}-d_{1}^{2}+4d_{1}+d_{2}d_{3}^{2}+d_{3}d_{4}^{2}+14&d=(d_{1},d_{3},d_{4},d_{5},d_{2}).\end{cases}$
Case 2:
In this case, randomly picking vertex values and considering the elimination of matched results, so we have
$\sigma(T_{2})=\begin{cases}2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{1}d_{2}^{2}+d_{3}^{2}+d_{4}^{3}+10&d=(d_{5},d_{1},d_{2},d_{3},d_{4}),\\ 2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}d_{3}^{2}+d_{2}^{3}+d_{4}^{3}+21&d=(d_{5},d_{1},d_{4},d_{2},d_{3}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{2}d_{3}^{2}+d_{3}d_{4}^{2}+d_{2}^{3}+22&d=(d_{1},d_{5},d_{2},d_{4},d_{3}),\\ 2d_{1}^{3}-4d_{1}^{2}+5d_{1}+d_{1}d_{2}^{2}+d_{2}d_{3}^{2}+d_{4}^{3}+8&d=(d_{5},d_{4},d_{1},d_{3},d_{2}),\\ 2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}d_{3}^{2}+d_{2}^{3}+d_{4}^{3}+24&d=(d_{5},d_{2},d_{4},d_{1},d_{3}).\end{cases}$
Case 3:
In this case, To analyse these cases, we will debate each case with the others and then extract the maximum case value and demonstrate it through the following discussion as:
$\sigma(T_{3})=\begin{cases}d_{1}^{3}+2d_{1}^{2}+d_{1}+d_{2}d_{3}^{2}-d_{1}d_{2}^{2}+d_{4}^{3}-d_{3}d_{4}^{2}-d_{2}^{3},&d_{1}\geqslant 2\\ d_{1}^{3}-3d_{1}^{2}-d_{1}+d_{1}d_{2}^{2}-d_{2}d_{3}^{2}+d_{2}^{3}-11,&d_{1}\geqslant 3\\ d_{1}^{2}-2d_{1}+d_{3}d_{4}^{2}-d_{3}^{2}+d_{3}^{3}-d_{4}^{3}+4&d_{1}\geqslant 3\\ d_{1}^{2}-2d_{1}+d_{3}d_{4}^{2}-d_{3}^{2}+d_{3}^{3}-d_{4}^{3}+9&d_{1}\geqslant 2\\ d_{1}^{3}+4d_{1}^{2}-d_{1}+d_{2}d_{3}^{2}-d_{1}d_{2}^{2}+d_{3}d_{4}^{2}-d_{3}^{2}-d_{4}^{3}+4&d_{1}\geqslant 3\\ d_{1}d_{2}^{2}-d_{2}d_{3}^{2}+d_{3}^{2}-d_{2}^{3}-14&d_{1}\geqslant 3\\ -d_{1}^{2}+2d_{1}-d_{3}d_{4}^{2}+d_{4}^{3}-1&d_{1}\geqslant 5\\ -2d_{1}+d_{3}d_{4}^{2}-d_{1}d_{2}^{2}+d_{2}^{3}-d_{4}^{3}+14&d_{1}\geqslant 3.\end{cases}$
Notice that from all cases,there are many objectionable relationships because they do not give the required results according to the terms we have established, so they were excluded and the cases that meet these terms were discussed, through which we obtain, so that we have:
$\sigma(T)=\begin{cases}\sigma_{\max}(T)=2d_{1}^{3}-4d_{1}^{2}+5d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}+d_{4}^{3}+8\\ \sigma_{\min}(T)=2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{2}d_{3}^{2}+d_{3}^{3}+19.\end{cases}$
As desire. ∎ In Table 1 we show some values of degree sequence according to Hypothesize 2 for analyse these cases with updated calculations the modified formulas. Each row pairs vertices differing only in $d_{5}$ as we show that.
Table 1. Values of $\sigma_{\max}(T)$ and $\sigma_{\min}(T)$ for Selected Parameters
$(d_{1},\ldots,d_{5})$ $\sigma_{\max}(T)$ $\sigma_{\min}(T)$ $(d_{1},\ldots,d_{5})$ $\sigma_{\max}(T)$ $\sigma_{\min}(T)$
(4,7,12,15,20) 6363 3027 (4,7,12,15,23) 6363 3027
(4,7,12,17,20) 8669 3027 (4,7,12,17,23) 8669 3027
(4,7,14,15,20) 7887 5955 (4,7,14,15,23) 7887 5955
(4,10,12,17,20) 8873 3663 (4,10,12,17,23) 8873 3663
(4,10,14,15,20) 8091 6591 (4,10,14,15,23) 8091 6591
(4,10,14,17,20) 10397 6591 (4,10,14,17,23) 10397 6591
(6,7,12,15,20) 9885 5629 (6,7,12,15,23) 9885 5629
(6,7,12,17,20) 13191 5629 (6,7,12,17,23) 13191 5629
(6,7,14,15,20) 11409 8557 (6,7,14,15,23) 11409 8557
(6,7,14,17,20) 14715 8557 (6,7,14,17,23) 14715 8557
(6,10,14,15,20) 11613 9193 (6,10,14,15,23) 11613 9193
(6,10,14,17,20) 14919 9193 (6,10,14,17,23) 14919 9193
We want to refer to Hypothesize 2 according to the constant term $(d_{1}-1)^{3}+(d_{4})^{3}$ that is employed in Hypothesize 3 strongly influence outcomes for amplifying the term $(d_{1}-1)^{3}+(d_{6}-1)^{3}$.
Hypothesize 3.
Let be a sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then Sigma index is:
$\sigma(T)=\sum_{i=1}^{n-2}d_{i}(d_{i+1})^{2}+(d_{1}-1)^{3}+(d_{6}-1)^{3}.$
where inequality holds if and only if $d_{i}=d_{i-1}+1$.
Proof.
We will discuss two basic conditions in the proof, one of which leads to the other, as it is fulfilled under certain conditions as follows:
Case a:
In this case we assume the sequence is $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, in this case the term is $d_{i}=d_{i-1}+1$, the number of possible cases in this case is 720 cases, some of which give similar results and some of which do not fulfil the condition and give negative values, so we took 10 cases and compared them with the rest of the cases and reviewed the results of this comparison through cases from case 1 to case 10.
Case b:
In this case we assume the sequence is $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, this case is a generalisation of the previous case but with no specific conditions, just a normal order.
Therefore, let begin with first Case a as following:
Case 1:
in this case we have
$\sigma(T_{1})=\begin{cases}2d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{1}d_{2}^{2}+d_{2}d_{3}^{2}+d_{3}d_{4}^{2}+d_{5}^{3}+4&d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6}),\\ 2d_{1}^{3}+(d_{2})(d_{3})^{2}+2d_{3}(d_{4})^{2}+d_{1}(d_{2})^{2}-3d_{1}^{2}+3d_{1}+33&d=(d_{2},d_{4},d_{5},d_{3},d_{6},d_{1}),\\ 2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}d_{3}^{2}+d_{1}d_{2}^{2}+d_{3}^{3}+d_{5}^{3}+24&d=(d_{4},d_{5},d_{3},d_{1},d_{2},d_{6}),\\ 2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}^{3}+d_{3}^{3}+d_{3}d_{4}^{2}+d_{4}d_{5}^{2}+54&d=(d_{3},d_{1},d_{6},d_{2},d_{5},d_{4}),\\ 2d_{1}^{3}-5d_{1}^{2}+5d_{1}+d_{2}^{3}+d_{2}d_{3}^{2}+d_{3}d_{4}^{2}+d_{4}^{3}+17&d=(d_{3},d_{1},d_{2},d_{5},d_{4},d_{6}).\end{cases}$
Case 2:
in this case we have the constant term is differ $d_{1}$, by considering $\beta=d_{2}d_{3}^{2}+d_{3}^{3}+2d_{3}^{2}+d_{3}+d_{5}^{3}$, $\gamma=d_{4}d_{5}^{2}+2d_{4}^{2}+d_{2}d_{4}^{2}$ and $\zeta=d_{1}d_{2}^{2}+d_{3}d_{5}^{2}+d_{5}^{2}$ then we have:
$\sigma(T_{2})=\begin{cases}3d_{1}^{3}-3d_{1}^{2}+6d_{1}+d_{3}d_{4}^{2}+d_{3}^{3}+d_{5}^{3}+17&d=(d_{6},d_{5},d_{3},d_{2},d_{1},d_{4}),\\ (d_{2})^{3}+(d_{1}-1)d_{1}^{2}+(d_{1}-2)(d_{1}-1)^{2}+\beta+13&d=(d_{3},d_{2},d_{1},d_{4},d_{5},d_{6}),\\ d_{5}^{3}+3d_{1}^{3}-2d_{1}^{2}+4d_{1}+d_{1}d_{3}^{2}+d_{3}^{2}+d_{4}^{3}+19&d=(d_{6},d_{3},d_{2},d_{1},d_{4},d_{5}),\\ 2d_{1}^{3}-4d_{1}^{2}+3d_{1}+2d_{2}^{3}+2d_{2}^{2}+d_{2}+\gamma+49&d=(d_{1},d_{6},d_{2},d_{4},d_{5},d_{3}),\\ 2d_{1}^{3}-3d_{1}^{2}+3d_{1}+2d_{3}^{3}+2d_{3}^{2}+d_{3}+\zeta+26&d=(d_{2},d_{1},d_{5},d_{6},d_{3},d_{4}).\end{cases}$
Through these discussed cases, we can observe that both elements at the vertices of the tree have a constant coefficient of $(d_{i})^{3}$, and we also observe that the elements located in the centre have a coefficient of the form $d_{i}(d_{i+1})^{2}$, so the difference remains in the value of the constant added to these elements, which varies from one case to another, according to Case 1 we have:
$\sigma(T_{3})=\begin{cases}2d_{1}^{3}-3d_{1}^{2}+2d_{1}+d_{1}d_{2}^{2}+d_{3}d_{4}^{2}-d_{3}^{3}-2d_{3}^{2}-d_{3}-9,\\ d_{3}^{3}-d_{2}d_{3}^{2}+d_{4}d_{5}^{2}-d_{4}^{3}+37.\end{cases}$
According to Case 2 we have:
$\sigma(T_{4})=\begin{cases}3d_{1}^{3}-2d_{1}^{2}+2d_{1}+d_{3}d_{4}^{2}-d_{2}^{3}-d_{2}d_{3}^{2}-2d_{3}^{2}-d_{3}+4,\\ -d_{1}^{3}-4d_{1}^{2}+3d_{1}+d_{2}^{3}+d_{2}d_{3}^{2}-d_{3}^{3}-d_{1}d_{2}^{2}-d_{3}d_{5}^{2}-d_{5}^{2}+d_{5}^{3}-13.\end{cases}$
so that, notice terms dominate due to rapid growth according to $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, and constants value, then we have:
$\sigma(T)=\begin{cases}\sigma_{\max}(T)=2d_{1}^{3}-4d_{1}^{2}+3d_{1}+2d_{2}^{3}+2d_{2}^{2}+d_{2}+\gamma+49,\\ \sigma_{\min}(T)=(d_{2})^{3}+(d_{1}-1)d_{1}^{2}+(d_{1}-2)(d_{1}-1)^{2}+\beta+13.\end{cases}$
Now, let be considered the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$, then we have according Case b where considering $k>0,k\in\mathbb{N}$,and the number of pendent vertices $n\geq 1$ then:
Case 1:
In this case, we have the sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$, then we have:
$\sigma(T)$ $=(n-1)((d_{6}-1)^{2}+(d_{1}-1)^{2})+(n-2)((d_{2}-1)^{2}+$
$+(d_{3}-1)^{2}+(d_{4}-1)^{2}+(d_{5}-1)^{2})+k.$
$=(n-1)((d_{1}-1)^{2}+(d_{6}-1)^{2})+(n-2)(\sum_{i=2}^{5}(d_{i}-1)^{2})+k.$
Case 2:
In this case, we have the sequence $d=(d_{6},d_{4},d_{2},d_{1},d_{3},d_{5})$, then:
$\sigma(T)=(n-1)((d_{6}-1)^{2}+(d_{5}-1)^{2})+(n-2)((d_{4}-1)^{2}+(d_{2}-1)^{2}+(d_{1}-1)^{2}+(d_{3}-1)^{2})+k.$
Therefore, we have:
$\sigma(T)=(n-1)(\sum_{i=5}^{6}(d_{i}-1)^{2})+(n-2)(\sum_{i=2}^{4}(d_{i}-1)^{2})+k.$
Note that in both cases we need to know the bifurcations generated by each of the vertices, which we were able to do in Case a discussed above, so we can formulate anchors the maximum, while the smallest $d_{6}$ is absent in most expressions, reducing exact numerical bounds, specific values impact show in the following:
$\sigma(T)=\begin{cases}\sigma_{\max}(T)=(n-1)(\sum_{i=5}^{6}(d_{i}-1)^{2})+(n-2)(\sum_{i=2}^{4}(d_{i}-1)^{2})+k\\ \sigma_{\min}(T)=(n-1)((d_{1}-1)^{2}+(d_{6}-1)^{2})+(n-2)(\sum_{i=2}^{5}(d_{i}-1)^{2})+k.\end{cases}$
As desire. ∎
Theorem 3.9.
Let be $T$ a tree of order $n$ define a sequence $d=(d_{1},\dots,d_{n})$ where $d_{1}\geqslant d_{2}\geqslant\dots\geqslant d_{n}$, then Albertson index among tree $T$ is:
$\sigma(T)=(d_{n}-1)^{3}+(d_{1}-1)^{3}+\sum_{i=1}^{n}(d_{i}-d_{i+1})^{2}+\sum_{i=2}^{n-1}(d_{i}-1)^{2}(d_{i}-2).$
Theorem 3.10.
Let $T$ be a tree of order $n>0$, define degree sequence $d=(d_{1},\dots,d_{n})$ where $d_{n}\geqslant\dots\geqslant d_{1}$, then Sigma index among tree given as:
$\sigma(T)=\sum_{i\in\{1,n\}}(d_{i}+1)(d_{i}-1)^{2}+\sum_{i=2}^{n-1}(d_{i}+2)(d_{i}-1)^{2}+\sum_{i=2}^{n-1}(d_{i}-d_{i+1})^{2}+2n-2.$
Example 1.
Let be a degree sequence is $d=(4,8,10,14,18,20)$, then for Max sigma we have:
$[14802,18,4,10,14,8,20],[14802,18,4,10,14,8,20],[14802,18,4,14,8,10,20],[14802,18,4,14,10,8,20],$
$[14802,18,8,10,14,4,20],[14802,18,8,14,4,10,20],[14802,18,8,14,10,4,20],[14802,18,10,4,14,8,20],$
$[14802,18,10,8,14,4,20],[14802,20,4,10,14,8,18],[14802,20,4,14,8,10,18],[14802,20,4,14,10,8,18],$
$[14802,20,8,10,14,4,18],[14802,20,8,14,4,10,18],[14802,20,8,14,10,4,18],[14802,20,10,4,14,8,18],$
$[14802,20,10,8,14,4,18].$
and for Min sigma we have:
$[14196,4,10,14,18,20,8],[14196,4,10,14,18,20,8],[14196,4,10,14,20,18,8],[14196,4,10,18,20,14,8],$
$[14196,4,10,20,18,14,8],[14196,4,14,18,20,10,8],[14196,4,14,20,18,10,8],[14196,4,18,20,14,10,8],$
$[14196,4,20,18,14,10,8],[14196,8,10,14,18,20,4],[14196,8,10,14,20,18,4],[14196,8,10,18,20,14,4],$
$[14196,8,10,20,18,14,4],[14196,8,14,18,20,10,4],[14196,8,14,20,18,10,4],[14196,8,18,20,14,10,4],$
$[14196,8,20,18,14,10,4].$
Note the location of the vertices in the sequence relative to the maximum value and the minimum value as a result of considering all the possibilities for this relationship as we show that in Theorem 3.10.
4. Conclusion
Through this paper, we presented a study of the topological indices defined by the Sigma Index and the Albertson Index according to a degree sequence $d_{n}\geqslant\dots\geqslant d_{1}$, in Hypothesize 3 for a degree sequence $d=(d_{1},d_{2},d_{3},d_{4},d_{5},d_{6})$ where $d_{6}\geqslant d_{5}\geqslant d_{4}\geqslant d_{3}\geqslant d_{2}\geqslant d_{1}$, then we defined Sigma index as $\sigma(T)=\sum_{i=1}^{n-2}d_{i}(d_{i+1})^{2}+(d_{1}-1)^{3}+(d_{6}-1)^{3}$. The examination of topological indices on trees gives us a general overview through bounds to find the maximum and minimum bounds and this is what we were keen to show in the proofs we presented for each case discussed.
Acknowledgments
I would like to extend my deepest gratitude to my thesis advisor, Dr. Artem Kurnosov, for her invaluable guidance and support throughout this research project. I also appreciate the insightful feedback provided by Prof.Belov Alexey Yakovlevich during the review process.